### v1

# Dark Energy from Time Crystals

Abstract
In this work, we analyze a scalar field model which gives rise stable bound states in field space characterized by nonzero motion that breaks the underlying time translation symmetry of its Hamiltonian, known as time crystals. We demonstrate that an ideal fluid made up of these time crystals behaves as phantom dark energy characterized by an equation of state $w<-1$, speed of sound squared $c_{s}^{2}\geq 0$, and nonnegative energy density $\rho\geq 0$.
I Introduction
Non-canonical scalar field models have been extensively studied in cosmology due to their rich dynamics and potential to explain dark energy and inflationary phenomena [1, 2, 3, 5]. These models exhibit interesting behaviors that besides phantom regimes in cosmology [5], lead to non-trivial stability properties in condensed matter systems  [7, 6].
A representative class of models with non-canonical kinetic terms give rise to the formation of crystalline type structures which oscillate in time rather than space, known as time crystals. They arise whenever a system described by a time- independent Hamiltonian has stable bound state solutions that spontaneously break the time translation symmetry of its Hamiltonian and reduce its time translation symmetry to a discreet subgroup that is periodic in time.
In this work, we investigate the time crystals model proposed in  [6, 7] and show that they behave as a ideal fluid with a phantom dark energy equation of state. Let’s start our investigation with a general non-canonical Hamiltonian in order to derive expressions for the energy density $\rho$, pressure $p$, equation-of-state parameter $w$, and speed of sound squared $c_{s}^{2}$ and then specialize our results to the time crystal model of  [6].
A general non-canonical scalar field model is described by the Lagrangian density
$\mathcal{L}=f(\phi)g(X)-V(\phi),$ (1)
where: $\phi$ is a scalar field and dot denotes the time derivative; $X=\dfrac{1}{2}\dot{\phi}^{2}>0$; $f(\phi)>0$ is a positive semi-definite function of $\phi$; $g(X)$ is a non trivial function of $X$, and $V(\phi)$ is the potential of the scalar field which we take to be a semi-definite positive function.
Below we first inspect the conditions required for an ideal fluid with a phantom equation of state $w<-1$, a nonnegative energy density $\rho\geq 0$, and a positive speed of sound squared $c_{s}^{2}\geq 0$. Then, we apply these considerations to the specific model of time crystals of  [6, 7] and explore its connection to phantom dark energy.
II General Formalism
The energy-momentum tensor for a scalar field in a flat Friedmann-Robertson-Walker (FRW) universe is given by:
$T_{\mu\nu}=\partial_{\mu}\phi\dfrac{\partial\mathcal{L}}{\partial(\partial^{\nu}\phi)}-g_{\mu\nu}\mathcal{L}.$ (2)
For a homogeneous scalar field $\phi=\phi(t)$, the corresponding energy density $\rho$ and pressure $p$ of the ideal fluid are:
[…]
$p$ $=-\dfrac{1}{3}T^{i}_{\ i}=\mathcal{L},$ (4)
where $i$ runs over spatial indices.
The relevant Einstein equations that are sufficient for describing the ideal fluid are the Friedmann equation and the Bianchi identity. The Friedman equation is
$(\frac{\dot{a}}{a})^{2}=\frac{8\pi G}{3}(\rho+\rho_{m})$ (5)
where $a(t)$ is the scale factor of the FRW metric, ˙$\dot{}=d/dt$. In general, we allow for other matter and energy components $\rho_{m}$ in the universe, not related to the non-canonical ideal fluid $\phi$. However, the phantom dark energy component soon dominates the expansion while the matter and radiation components dilute away, rendering them insignificant. For this reason we ignore the contribution of $\rho_{m}$ in the equations below.
The Bianchi identity of the fluid is
$\dot{\rho}+3\frac{\dot{a}}{a}(\rho+p)=0$ (6)
In this class of models, the non canonical conjugate momenta $\Pi$, obtained as a derivative of $\mathcal{L}$ with respect to $\dot{\phi}$, is different from the kinetic momentum $\dot{\phi}$
$\Pi=\dfrac{\partial\mathcal{L}}{\partial\dot{\phi}}=f(\phi)g^{\prime}(X)\dfrac{\partial X}{\partial\dot{\phi}}=f(\phi)g^{\prime}(X)\dot{\phi},$ (7)
where $g^{\prime}(X)=\dfrac{dg}{dX}$ and $\dot{\phi}=\pm\sqrt{2X}$.
For an ideal fluid the Hamiltonian of the scalar field serves as the energy density of the cosmic fluid and the Lagrangian as the pressure term, Eqns. (4). Due to the non-canonical nature of the Lagrangian, the Lagrangian and the Hamiltonian expressions below are quite different from one another. Substituting this into Eq. (3) to obtain the Hamiltonian
$\rho$ $=\mathcal{H}=\Pi\dot{\phi}-\mathcal{L}=\dot{\phi}\left(f(\phi)g^{\prime}(X)\dot{\phi}\right)-\left(f(\phi)g(X)-V(\phi)\right)$
$=f(\phi)\left(2Xg^{\prime}(X)-g(X)\right)+V(\phi).$ (8)
whereas from Eq. (4), we have:
$p=f(\phi)g(X)-V(\phi).$ (9)
The equation-of-state parameter $w$ of the cosmic fluid $\phi$ is defined by:
$w_{\phi}=\dfrac{p}{\rho}=\dfrac{f(\phi)g(X)-V(\phi)}{f(\phi)\left(2Xg^{\prime}(X)-g(X)\right)+V(\phi)}.$ (10)
The speed of sound squared is given by [4]:
$c_{s}^{2}=\dfrac{\partial p}{\partial\rho}=\dfrac{p_{,X}}{\rho_{,X}},$ (11)
where $p_{,X}=\dfrac{\partial p}{\partial X}$ and $\rho_{,X}=\dfrac{\partial\rho}{\partial X}$.
Computing $p_{,X}$ and $\rho_{,X}$ gives:
$p_{,X}$ $=f(\phi)g^{\prime}(X),$ (12)
$\rho_{,X}$ $=f(\phi)\left(g^{\prime}(X)+2Xg^{\prime\prime}(X)\right),$ (13)
with $g^{\prime\prime}(X)=\dfrac{d^{2}g}{dX^{2}}$.
Therefore, the expression for the speed of sound squared is
$c_{s}^{2}=\dfrac{g^{\prime}(X)}{g^{\prime}(X)+2Xg^{\prime\prime}(X)}.$ (14)
II.1 Conditions for Stable Phantom Behavior with Positive Energy density
A phantom dark energy fluid has an equation-of-state parameter $w_{\phi}<-1$.
From Eq. (10), the condition $w_{\phi}<-1$ requires
$\rho+p<0.$ (15)
where $\rho+p$ using Eqs. (8) and (9)
$\rho+p$ $=2f(\phi)Xg^{\prime}(X).$ (16)
The condition for $w_{\phi}<-1$ then implies
$f(\phi)Xg^{\prime}(X)<0.$ (17)
Since $f(\phi)>0$ and $X>0$, the condition for a phantom dark energy equation of state further simplifies to
$g^{\prime}(X)<0.$ (18)
Note that the conditions for phantom behavior depends solely on the expression of the kinetic term $g(X)$ and are completely independent of the type of potential $V(\phi)$ chosen.
We further demand that the energy density of the ideal phantom dark energy fluid given by Eqn. (8) is non-negative. The requirement $\rho\geq 0$ yields
$f(\phi)\left(2Xg^{\prime}(X)-g(X)\right)+V(\phi)\geq 0.$ (19)
Since $f(\phi)>0$, we can further simplify it as:
$2Xg^{\prime}(X)-g(X)+\dfrac{V(\phi)}{f(\phi)}\geq 0$ (20)
This condition depends on the specific forms of $g(X)$, $V(\phi)$, and $f(\phi)$. However, given that $g^{\prime}(X)\leq 0$ for phantom dark energy ideal fluids, and further restrictions that we will discuss below in the context of time crystals, make the constraints on the potential energy term sufficiently weak to be accommodated by a large class of generic models. Potentials which satisfy the positive energy condition would only be weakly tuned.
If the speed of sound squared were negative we would be looking at a tachyonic fluid which can be unstable. Therefore, requiring the speed of sound squared of the fluid to be positive or zero provides the regions of stability in the field space.
The speed of sound squared is given by Eq. (14). The requirement of $c_{s}^{2}\geq 0$ leads to the following conditions on derivatives of $g(X)$
$g^{\prime}(X)\left[g^{\prime}(X)+2Xg^{\prime\prime}(X)\right]\geq 0.$ (21)
This inequality depends on the signs of $g^{\prime}(X)$ and $g^{\prime}(X)+2Xg^{\prime\prime}(X)$. But the equation of state of the phantom fluid already imposed the condition that $g^{\prime}(X)\leq 0$ , which means that $c_{s}^{2}\geq 0$ can only be satisfied for
$2Xg^{\prime\prime}(X)+g^{\prime}(X)\leq 0$ (22)
Combining the above conditions ($c_{s}^{2}\geq 0$, $w<-1$, $\rho\geq 0$) for the phantom regime, we now compare these results against the time crystal model given in [6]. We show that the time crystal model of  [6, 7] gives rise to phantom dark energy and we seek to identify the ranges of $X$ where the phantom behavior is stable and the energy density is non-negative. This requires solving the inequalities described in this section simultaneously.
III Application of Phantom Dark Energy Conditions to the Time Crystal Model
Time crystals were first discovered in [7, 6]. Although their discovery is relatively recent, there already exists a vast literature on the topic of time crystals and experiments set up for observing them in a lab, reviewed in [8].
In this work, we take the kinetic term $g(X)$ to be that of the time crystal model of [6] and show that it produces phantom dark energy. However, the formalism presented here can be extended to more complex time crystal models.
III.1 A Time Crystals Lagrangian
Consider the Lagrangian density,
$\mathcal{L}=-\kappa_{2}X+\lambda_{2}X^{2}-V(\phi),$ (23)
proposed for time crystals in [6] where $\kappa_{2}>0$ and $\lambda_{2}>0$ are constants, $f(\phi)=1$, and $V(\phi)$ is a positive semi definite potential. In what follows, the non canonical kinetic term $g(X)$ of the previous section is therefore chosen to be that of the time crystal
$g(X)=-\kappa_{2}X+\lambda_{2}X^{2}$ (24)
Applying the results of the previous section, the Hamiltonian obtained from this Lagrangian, is
$\mathcal{H}=-\kappa_{2}X+3\lambda_{2}X^{2}+V(\phi)$ (25)
which can be rewritten as
$\mathcal{H}=3\lambda_{2}(X-X_{t})^{2}+(V(\phi)-\frac{\kappa_{2}^{2}}{12\lambda_{2}})$ (26)
where
$\mathcal{H}+\mathcal{L}=\Pi\dot{\phi}$ (27)
and conjugate momentum,
$\Pi=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=\dot{\phi}(-\kappa_{2}+2\lambda_{2}X)$ (28)
Based on these relations and Lagrangian equations, the meaning of $X_{t}=\kappa_{2}/6\lambda_{2}$ will soon become clear to define the turning points of the bound orbit of the time crystal that minimize energy, corresponding to $\dot{\phi}_{t}=\pm(\kappa_{2}/3\lambda_{2})^{1/2}\neq 0$. Importantly, there is motion $X\neq 0$ along the orbit $X_{t}$ which gives rise to the spontaneous breaking of the time translation symmetry of the Hamiltonian as is apparent from Eqn. (26). The effective potential $W(\phi)$ in the expression for the Hamiltonian, Eqn. (26), is given by
$W(\phi)=V(\phi)-\kappa_{2}^{2}/12\lambda_{2}=V(\phi)-\Lambda_{T}$ (29)
Interestingly, note that there is an overall constant negative energy term $\Lambda_{T}$ in Eqn. (29) originating from time crystals contributing to the effective potential of the energy expression of Eqn. (26). The term $\Lambda_{T}$ which appears naturally from time crystals, can offset a large positive vacuum energy and alleviate the notorious fine tuning of problem of vacuum energy (by $123-$ orders of magnitude), through a far less tuned choice of its parameters $\lambda_{2},\kappa_{2}$. The parameters $\lambda_{2},\kappa_{2}$ which enter in a ratio, can take reasonable values and still be such that their combination in the time crystal contribution to vacuum energy $\Lambda_{T}$ is large or small enough to absorb the background’s positive vacuum energy.
Furthermore, we can also rewrite the Lagrangian for this model as
$\mathcal{L}=\lambda_{2}(X-X_{d})^{2}-(V(\phi)+\frac{\kappa_{2}^{2}}{4\lambda_{2}})$ (30)
As we will see below, by writing the Hamiltonian and Lagrangian in this manner, both points $X_{t},X_{d}$ will play an important role in the dynamics of the fluid.
For the ideal fluid with a stress energy tensor,
$T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}-pg_{\mu\nu}$ (31)
with $u_{\mu}$ the unit 4-vector normalized to unity, whereas the pressure density of the fluid is given by the Lagrangian Eq. (30) of the scalar field, $p=\mathcal{L}$, and the energy density of the fluid is given by the Hamiltonian expression $\rho=\mathcal{H}$ of Eq. (26). With this notation, we can rewrite the energy density of the fluid:
$\rho=3\lambda_{2}(X-X_{t})^{2}+W(\phi)$ (32)
Note that while the non-canonical form of the kinetic term of the time crystal contributes a term $\Lambda_{t}$ which reduces the total energy of the ideal fluid by a finite negative constant energy $-\Lambda_{T}=-\frac{\kappa_{2}^{2}}{12\lambda_{2}}$, meanwhile the correction to the pressure of the ideal fluid increases the pressure by a positive constant amount $p_{t}=3\Lambda_{T}=+\kappa_{2}^{2}/4\lambda_{2}$. The pressure correction term for the effective pressure is positive and three times larger than the energy correction in the effective potential, typical of a solids equation of state.
From Lagrangian equations of motion we have
$\frac{\partial\mathcal{L}}{\partial\phi}-3H\Pi=\frac{d}{dt}\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=\frac{d}{dt}\Pi=\dot{\Pi}$ (33)
where $H=\frac{\dot{a}}{a}$ is the Hubble parameter.
These equations of motion are directly related to three points and regions of interest in the field space of the time crystals ideal fluid, that show up as $X_{t}$ and $X_{d}$ in Eqns.  (26) and  (30).
The requirement that kinetic energy is minimized leads to
$\dot{\Pi}=\ddot{\phi}(-\kappa_{2}+3\lambda_{2}\dot{\phi}^{2})=0$ (34)
which, for the nontrivial bound state $\dot{\phi}\neq 0$, implies that $\Pi(X)$ is a constant of motion along the turning points of time crystal orbit $\dot{\phi_{t}}=\pm\sqrt{\kappa_{2}/3\lambda_{2}}$ or $X=X_{t}$.
The turning points can be obtained by minimizing kinetic energy or equivalently from the cusps of the conjugate momentum. The cusps of the conjugate momentum $\Pi(X)$ indicate critical points in the fields space because the speed of sound diverges there $c_{s}^{2}->\infty$. these cusps are found by
$\frac{\partial\Pi}{\partial\dot{\phi}}=0$ (35)
solutions of which occur exactly at the turning points of the time crystal $\dot{\phi}_{t}=\pm\sqrt{\kappa_{2}/3\lambda{2}}$ or $X=X_{t}$. The latter, in combination with the fact that the speed of sound squared diverges at the boundary, indicate that these bound structures, the time crystals, are stable against perturbations and the field is confined to the time crystal region $X\leq X_{t}$. Hence, the time crystal bound state is a state of minimum energy, despite that the underlying time translation symmetry is broken and there is motion $X\neq 0$ along the orbit. The meaning of ($X_{t},\pm\phi_{t}$) then becomes clear as the turning points that define the orbit of the bound state, but which, due to motion break time translation symmetry of the Hamiltonian and become time crystals.
Another point of interest is the one where the group velocity of the field vanishes and the kinetic energy is once again minimized, the point where $\Pi=\frac{\mathcal{L}}{\partial\dot{\phi}}=0$. As can be seen from Eq. (28), this occurs at $X=X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ or equivalently along the orbit defined by the special points $\phi_{d}=\pm\sqrt{\kappa_{2}/\lambda_{2}}$ that enter the Lagrangian expression, Eqn. (30).
III.2 Phantom Dark Energy from the Time Crystal model
Let us now investigate the relation between the time crystal model described in the previous section to phantom dark energy fluids with ($w<-1$, $c_{s}^{2}>0$, $\rho>0)$. The first and second derivatives of $g(X)$ are:
$g^{\prime}(X)$ $=-\kappa_{2}+2\lambda_{2}X,$ (36)
$g^{\prime\prime}(X)$ $=2\lambda_{2}.$ (37)
The condition for Phantom dark energy behavior, ($w<-1$), from Eq. (18) is: $g^{\prime}(X)<0$. Substituting $g^{\prime}(X)$ from Eq. (36):
$-\kappa_{2}+2\lambda_{2}X<0.$ (38)
and solving for $X$:
$X<\dfrac{\kappa_{2}}{2\lambda_{2}}.$ (39)
Hence, phantom behavior occurs anywhere in the region below the point $X=X_{d}$ where group velocity of the field and conjugate momentum $\Pi$ vanish:
$0<X<\dfrac{\kappa_{2}}{2\lambda_{2}}.$ (40)
The condition for nonnegative energy density ($\rho\geq 0$), with $f(\phi)=1$ is given by:
$-\kappa_{2}X+3\lambda_{2}X^{2}+V(\phi)\geq 0.$ (41)
For potentials that satisfy $W(\phi)\geq 0$, equivalently $V(\phi)\geq\frac{\kappa_{2}^{2}}{12\lambda_{2}}$, the energy density is positive everywhere, as per Eqn.  (26).
The general case of $V(\phi)$ in the above equation is a quadratic inequality in $X$ which can be solved for a range of $X$ where $\rho\geq 0$.
Given the bounds on $g(X),W(\phi)$ it is straightforward to check that the energy density is always positive in the region of interest where the ideal fluid as phantom behavior, which included the time crystal $X\leq X_{t}$.
The equation of state parameter $w_{\phi}$ for the ideal fluid described by the time crystals model is given by
$w_{\phi}=\frac{p}{\rho}=\frac{-\kappa_{2}X+\lambda_{2}X^{2}-V(\phi)}{-\kappa_{2}X+3\lambda_{2}X^{2}+V(\phi)}\leq-1$ (42)
for $X\leq X_{d}$.
Finally let’s check the condition for stability ($c_{s}^{2}\geq 0$) of the phantom fluid.
From Eq. (14), we have:
$c_{s}^{2}=\dfrac{-\kappa_{2}+2\lambda_{2}X}{-\kappa_{2}+6\lambda_{2}X}.$ (43)
The requirement $c_{s}^{2}\geq 0$, implies:
$\left(-\kappa_{2}+2\lambda_{2}X\right)\left(-\kappa_{2}+6\lambda_{2}X\right)\geq 0.$ (44)
Define:
$N$ $=-\kappa_{2}+2\lambda_{2}X,$ (45)
$D$ $=-\kappa_{2}+6\lambda_{2}X.$ (46)
Let’s consider the regions:
1. The time crystal region: $X\leq\dfrac{\kappa_{2}}{6\lambda_{2}}$
  • $N<0$,
  • $D<0$,
  • $c_{s}^{2}>0,w_{\phi}<-1,\rho>0$.
2. The forbidden region bounded below by the orbit of the time crystal $X_{t}$ and above by the zero group velocity $X_{d}$ respectively: $\dfrac{\kappa_{2}}{6\lambda_{2}}<X<\dfrac{\kappa_{2}}{2\lambda_{2}}$
  • $N<0$,
  • $D>0$,
[…]
3. The quintessence region: $X\geq X_{d}=\dfrac{\kappa_{2}}{2\lambda_{2}}$
  • $N\geq 0$,
  • $D\geq 0$,
  • $c_{s}^{2}\geq 0,w_{\phi}>-1,\rho>0$.
There is an additional point of interest in the forbidden region $\it{2}$, where the speed of sound goes through zero again and halfway in this region it switches from negative to positive indicating stability, at $X=X_{0}=\dfrac{\kappa_{2}}{4\lambda_{2}}=X_{d}/2$ found by the condition of Section $\it{1}$: $2Xg"(X)\leq-g^{\prime}(X)$. In the forbidden region $X_{t}\leq X\leq X_{0}$ the momentum $\Pi$ obtained by the quadratic equation above $\rho\geq 0$ is imaginary, again indicating that the field would be damped to zero quickly in this region. Therefore, stability ($c_{s}^{2}\geq 0$) occurs only when the field is confined in the time crystal region or when it is specially prepared to start at high conjugate moment for $X\geq X_{d}$:
$X\leq\dfrac{\kappa_{2}}{6\lambda_{2}}\quad\text{or}\quad X\geq\dfrac{\kappa_{2}}{2\lambda_{2}}.$ (47)
However, the unstable region above $X_{t}$ is inaccessible to the field since time crystals are in stable bound states and the speed of sound squared diverges at the boundary $X=X_{t}$, as given by cusps in the momentum space. These imply that the field is confined in the crystal region and it would require infinite work done to it to jump over the crystal boundary and to break the crystal structure. Even if the field were to tunnel through $X_{t}$ boundary with a small probability it would die out to zero quickly since the momentum becomes imaginary in the forbidden region. In other words, these crystals are very stable structures.
The only region accessible to the field is the time crystal interval that simultaneously satisfies all the conditions for the phantom dark energy which is the time crystal region: $0<X\leq\dfrac{\kappa_{2}}{6\lambda_{2}}$, provided $W(\phi)\geq 0$. In this region, we find the following.
$w_{\phi}<-1$ (phantom behavior), $c_{s}^{2}\geq 0$ (stability), $\rho\geq 0$ if the potential $V(\phi)$ is sufficiently large to ensure $\rho\geq 0$.The latter is automatically satisfied throughout the field space if we require $W(\phi)\geq 0$, which can be easily achieved by multiplying the potential $V(\phi)$ by an overall constant factor without changing or tuning its functional form.
For $X\geq\dfrac{\kappa_{2}}{2\lambda_{2}}$, stability is maintained and $\rho\geq 0$, but $w_{\phi}>-1$, so the phantom behavior is lost and replace by a quintessence type dynamics. However, if the field is bound below $X_{d}$ in the time crystal region it would not be possible for the field to break away from its bound state and suddenly jump over to $X=X_{d}$ which is the boundary of the quintessence region since the momentum $\Pi_{d}=0$ and the group velocity are zero at that initial boundary. The field could not access the region $X>X_{t}$.
Let’s classify the points and regions of interest in field space in the table below.
Based on the equation of motions for $X$, the behavior of $w_{\phi}$, $c^{2}_{s}$, and $\rho$ we showed that the field space is divided into distinct regions of interest, one of which contains the time crystals. Table 1 summarizes these results, detailing the range of $X$, energy density, equation-of-state parameter, and sound speed squared $c_{s}^{2}$ for all the field space regions.
Table 1: Classification of Key Points and Regions in the Time Crystal Scalar Field Model
Region Range of $X$ $\rho$ $w_{\phi}$ $c_{s}^{2}$
Trivial canonical vacuum energy $X=0$ $\rho=V(\phi)>0$ $w_{\phi}=-1$ $c_{s}^{2}=1$
Time Crystal Region $0<X<X_{t}=\frac{\kappa_{2}}{6\lambda_{2}}$ $\rho>0$ $w_{\phi}<-1$ $c_{s}^{2}>1$
Time Crystal Boundary $\lim_{X\rightarrow X_{t}^{-}=\frac{\kappa_{2}}{6\lambda_{2}}}$ $\rho=-\frac{\kappa_{2}^{2}}{12\lambda_{2}}+V(\phi)>0$ $w_{\phi}\rightarrow w_{\phi}^{\rm min}<-1$ $c_{s}^{2}\to\infty$
[…]
Boundary of quintessence $\Pi=0,X=X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ $\rho=\frac{\kappa_{2}^{2}}{4\lambda_{2}}+V(\phi)>0$ $w_{\phi}=-1$ $c_{s}^{2}=0$
quintessence $X>X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ $\rho=\frac{\kappa_{2}^{2}}{4\lambda_{2}}+V(\phi)>0$ $w_{\phi}>-1$ $10<c_{s}^{2}<$ 1
The Hamiltonian is a single-valued function at the trivial point $X=0$, the equation of state $w_{\phi}=-1$ is that of a pure positive cosmological constant and the speed of sound $c_{s}^{2}=1$ indicates a canonical ideal fluid.
In the region between zero and the turning points for the orbit $0<X<X_{t}=\frac{\kappa_{2}}{6\lambda_{2}}$ we have the time crystal which is a stable bound state, for the reasons discussed above. The Hamiltonian is a double-valued function of $\dot{\phi}$. The energy density remains positive:
$\rho_{\phi}=-\kappa_{2}X+3\lambda_{2}X^{2}+V(\phi)>0.$ (48)
The equation-of-state parameter satisfies:
$w_{\phi}<-1,$ (49)
characterizing phantom-like behavior. The sound speed squared is positive:
$c_{s}^{2}=\frac{-\kappa_{2}+2\lambda_{2}X}{-\kappa_{2}+6\lambda_{2}X}>0.$ (50)
This indicates stable perturbations within this interval. At the boundary $X=X_{t}$ the energy density reaches:
$\rho=-\frac{\kappa_{2}^{2}}{12\lambda_{2}}+V(\phi)>0.$ (51)
Here, the equation-of-state parameter $w_{\phi}$ attains its minimum value within the phantom regime ($w_{\phi}<-1$). The sound speed squared $c_{s}^{2}$ approaches infinity at the crystal’s boundary, indicating a stable time crystal structure.
The field would require an infinitely large pressure gradient, or force, to break out of the crystal boundary $X_{t}$. In a sense, this boundary acts as a hard wall potential for the forbidden region $X>X_{t}$, with the field trapped on its other side $X\leq X_{t}$. Hence, the region $X_{t}<X<X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ is inaccessible to the field. The speed of sound squared being greater than one is typical of crystalline structures. Its contribution to the metric perturbation equation is to dampens perturbations that would contribute to cold dark matter, thereby offering a smoking gun for testing this model.
In the region above the time crystal orbit, the scalar field continues to exhibit phantom-like behavior with
$w_{\phi}<-1.$ (52)
The energy density remains positive:
$\rho>0.$ (53)
However, the sound speed squared becomes negative and the momentum imaginary
$c_{s}^{2}=\frac{-\kappa_{2}+2\lambda_{2}X}{-\kappa_{2}+6\lambda_{2}X}<0,$ (54)
The speed of sound squared goes from negative values through zero halfway in the region at $X=X_{0}=X_{d}/2$ and then it switches to positive values.
At the next boundary $X=X_{d}$ where the conjugate momentum $\Pi$ and group velocity start from zero, the energy density is:
$\rho_{\phi}=\frac{\kappa_{2}^{2}}{4\lambda_{2}}+V(\phi)>0.$ (55)
The equation-of-state parameter returns to $-1$:
$w_{\phi}=-1,$ (56)
resembling a cosmological constant. The sound speed squared $c_{s}^{2}$ is unity:
$c_{s}^{2}=1.$ (57)
This ensures stable perturbations at this critical point.
Above $X_{d}$ in the region where $X_{d}<X$, the conjugate momentum $\Pi$ starts increasing and the scalar field transitions to quintessence-like behavior with:
$-1<w_{\phi}<0.$ (58)
The energy density remains positive:
$\rho_{\phi}>0.$ (59)
The sound speed squared satisfies:
$0<c_{s}^{2}<1,$ (60)
indicating stable perturbations within the quintessence regime. However, for the reasons discussed above, regions above the boundary of the time crystal with $X\ >X_{t}$ are not accessible unless the field is especially prepared to start at high momentum. Even then, the time crystal region $X\leq X_{t}$ is favored as the state that minimize energy.
IV Conclusions
In this work, we investigated the time crystal model proposed in  [6, 7] and showed that in the context of cosmology, it gives rise to a phantom dark energy ideal fluid, with an equation of state $w_{\phi}\leq-1$, a positive speed of sound squared $c_{s}^{2}\geq 0$, and positive energy density $\rho\geq 0$, which drives the expansion of the universe towards the Big Rip.
Furthermore, the speed of sound in the crystal region $X\leq X_{t}$ is such that $c_{s}^{2}\geq 1$ typical of crystalline structures. Observationally this sound speed would serve to dampen cold dark matter perturbation providing a way of testing this model.
The speed of sound squared diverges at the boundary of the time crystal in field space $X_{t}$ which implies that these are very stable structures bound by a hard wall in field space, such that the field would require vast amount of work to be done for it to break away from the crystal region. However, if for some reason, the field was prepared to start at high conjugate momentum or it tunneled through the hard wall, it would enter a forbidden region $X_{t}<X<X_{d}$ where the field is exponentially decaying, separating it from the quintessence region $X>X_{d}$. Yet, not only is the field confined to remain in the crystal region $X\leq X_{t}$ but dynamics favors it to naturally be found there since the crystal state minimizes energy.
Finally, the crystalline structure is time, which behaves as a phantom dark energy ideal fluid, naturally contribute an overall constant negative vacuum energy, $\Lambda_{T}$. Without the need of fine-tuning that is typical of the cosmological constant problem, this contribution to vacuum energy can offset a positive cosmological constant, with a reasonable, less-tuned, choice for the values of the parameters $(\lambda_{2},\kappa_{2})$ whose ratio determines the magnitude of $\Lambda_{T}$.
Acknowledgment: L. Mersini-Houghton is grateful to Y. J. Ng, O. Akarsy, E. Di Valentino for useful discussions, and to the Klingsberg foundation and the Bahnson fund for their support.